OpenStudy (anonymous):

What is the solution to the equation? -4q + 8 = -2q - 16 A. q = -12 B. q = -4 C. q = 4 D. q = 12

OpenStudy (kaylaprincess):

add 2q to -4q = -2q subtract 8 from - 16 = -8 -2q = -8 solve from there

syed98 (syedmohammed98):

-4q + 8 = -2q - 16 -4q + 2q = -16 -8 (shifting of numbers) -2q = -24 q = -24/-2 q=12 :)

OpenStudy (anonymous):

You're right^ I had already got it, thanks either way! :)

syed98 (syedmohammed98):

Your Welcome :)

OpenStudy (kaylaprincess):

( sorry for my mistake :(